OpenStudy (anonymous):

The average density of earth is taken to be 5.46 gcm-3.The diameter of Earth is 12.8 X10 to the power of 6. Calculate the mass of the Earth

OpenStudy (anonymous):

density=mass/volume volume(of sphere)=4/3*pi*r^3 mass=?

OpenStudy (anonymous):

volume*density should give u the answer

OpenStudy (anonymous):

first divide the diameter by two to get the radius then calculate the volume by 4/3\[\pi\](radius)*3

OpenStudy (anonymous):

Use the density formula and proceed.